OpenStudy (notamathgenius):

A photograph shows the 215 g golf club traveling at 55.0 m/s just before it strikes a 46 g golf at rest on a tee. After the collision, the club travels (in the same direction) at 42.0 m/s. Use the law of conservation of momentum to find the speed of the golf ball just after impact..... I need help not answers

OpenStudy (sparklestaraa):

OMG wait hold on let me ask my fiancé

OpenStudy (sparklestaraa):

can u help with math

OpenStudy (sparklestaraa):

physics for math :P

OpenStudy (notamathgenius):

sure :)

OpenStudy (sparklestaraa):

yayyy :) ok now we just wait for him to answer

OpenStudy (sparklestaraa):

hes a university student and i hate physics xd

OpenStudy (notamathgenius):

okie dokie

OpenStudy (sparklestaraa):

P = momentum P1 = P2 (conservation of momentum) 200*55 = 200*40 + 46*Vg Vg = 65.2 m/s i found it online b4 asking him- https://answers.yahoo.com/question/index?qid=20100302183242AAyDO9b

OpenStudy (notamathgenius):

WOO Thanks, tag me in the math and ill help

OpenStudy (sparklestaraa):

kk :D few Questions :P

OpenStudy (sparklestaraa):

u there i tagged u in